Wednesday 24 June 2009
This page lists relevant problems occurred on the SwissQuantum network.
December 18th, 2010: Power cut at CERN
Approximately at 1PM, there was a power cut at CERN. Both systems CERN-UNIGE and CERN-hepia were down until systems at CERN were powered on again (Monday December, 20th at 10AM).
July 27th, 2010: Wrong maintenance manipulation at hepia
Friday afternoon, on a bad manipulation during a maintenance (not related to the SwissQuantum project), all the systems in the servers’ room at hepia were without power. The SwissQuantum systems has been manually restarted Monday morning. The Web site remained down until Tuesday morning. All system are now up and running.
June 29th, 2010: Air conditioning problem at hepia
The air conditioning system had a broke down approximately at 1PM. The temperature in the servers’ room raised up to approximately 45°C! The air conditioning was repaired at around 5PM. It took several hours to lower down the temperature.
Due to the high temperature, Bob system (hepia-CERN link) automatically shut down the internal optical platform to avoid any damage to the detectors. The optical platform was manually restarted at around 8PM.
Alice system (hepia-UNIGE link) do not has any detector and continued to operate. The high temperature only reduced the secret key rate: due to the temperature, the system’s visibility slightly decreased and the QBER increased.
December 2nd, 2009: Power cut at CERN
Approximately at 2AM, there was a power cut at CERN. Both systems CERN-UNIGE and CERN-hepia were down until systems at CERN were powered on again (approximately at 10AM).
August 20th, 2009 - September 2nd, 2009: KeyServer problems
The new software deployed on the QKD that handles the connection between the QKD and the KeyServer was buggy and consumed a lot of computational power (CPU). This caused — on the QKDs between hepia and UNIGE — a system overload that is clearly visible on the Secret Key Rates graph. The problem did not appear on the other QKDs links since these systems require less computational power as they produce less keys.
The fix has been deployed remotely on September 1st, 2009. Due to a wrong manipulation during the deployment on a device located at hepia, the system required a manual restart. Since the site was closed during the deployment, the QKD link hepia-UNIGE has been non-operational during all night.
On September 2nd, 2009 all problems were solved and all devices are correctly working again.
June 24th, 2009: WDM deployment
WDM were deployed on the CERN-UNIGE link and the GBic on the QKD had to be replaced. Due to this replacement, the QKD required a restart. The QKD were non-operational for some hours.